the length of a rectangle is eight more than two times the width. Find the width if the perimeter of the rectangle is eaual to 88m.
Answer by Cromlix(4381)

Hi there,
Make the width = 'x'
Length = 8 + 2x
Perimeter = 2 x Width + 2 x Length
88 = 2(x) + 2(8 + 2x)
88 = 2x + 16 + 4x
Collect like terms:
6x = 88 - 16
6x = 72
x = 12
Width = 12 metres
Length = 32 metres
Hope this helps :-)